explain the concept of c.o.m. how ti changes before &and after the motion.How it changes before &after collisions.  how it is related to different bodies?

Grade:11

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear shanky

answer to all these question requires atleast 2 lecture for detail  answer .here i am just giving you the basic concept.

if any system made of diferent body of mass m1,m2,m3...... and at a distance from the reference point from where COM is measered is x1,x2,x3,,,, respectively  .so COM of the system about that point is defibed as

COM = (m1*x1+m2*x2+m3*x3 +......)/(m1 +m2 +m3) =(∑mixi)/∑mi

now lets consider mass m1 .m2.m3 ..... are moving with velocity V1,V2,V3.....respectively

then the COM is moving =(m1*v1+m2*v2+m3*v3 +......)/(m1 +m2 +m3) =(∑mivi)/∑mi

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin